40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Mar 6 15:24:05 2016, gautam, Update, CDS, FB down again 58.png
    Reply  Mon Mar 7 20:40:02 2016, ericq, Update, CDS, FB down again 
Message ID: 12024     Entry time: Sun Mar 6 15:24:05 2016     Reply to this: 12025
Author: gautam 
Type: Update 
Category: CDS 
Subject: FB down again 

I came in to check the status of the nitrogen and noticed that the striptool panels in the control room were all blank.

  • PMC was unlocked but I was able to relock it using the usual procedure
  • FB seems to be down: I was unable to ssh into it (or any of the FEs for that matter). I checked the lights on the RAID array, they are all green. I am holding off on doing a hard reboot of FB in case there is some other debugging that can be done first
  • None of the watchdogs were tripped, but judging by the green spots on the mirrors, all of them are moving quite a bit. I've shutdown the watchdogs on all the optics except the MC mirrors, but the ITMs and ETMs still seem to be moving quite a bit.

I am leaving things in this state for now. It is unclear why this should have happened, it doesn't seem like there was a power glitch?

Attachment 1: 58.png  3.451 MB  | Show | Show all
ELOG V3.1.3-